# Same Content, Different Representations: A Controlled Study for Table QA

### v2

@@ 4.3 Evaluation
Previous work (Deng et al., 2022; Wu et al., 2025a; Zhu et al., 2021) has largely relied on traditional metrics such as Exact Match (EM) and Partial Match (PM). However, these metrics suffer from well-known limitations: they depend on strict string-level comparison with the gold answer. Consequently, they penalize semantically correct predictions due to paraphrasing, formatting, or minor rounding differences. To address these limitations, we adopt LLMs as judges, a strategy shown to correlate strongly with human evaluation in a variety of reasoning tasks (Maekawa et al., 2025). Concretely, GPT-4o is given both the gold answer and the model’s prediction and prompted to determine semantic correctness. This evaluator tolerates surface variation while enforcing strict semantic equivalence. To assess the reliability of our LLM-based evaluator, we randomly sampled 100 evaluation cases across all datasets and model families. Each case was independently annotated by a human and compared to the LLM’s judgment. We observed a 96% agreement rate, demonstrating that the LLM-based evaluation is sufficiently stable for large-scale benchmarking. Unless stated otherwise, all results reported follow this evaluation protocol.
@@ 5.6 Case Study
We now present illustrative examples that capture the main failure modes when shifting from structured to semi-structured representations. Figure 7 shows a representative case from the “supplier” table, where the query asks for the top ten suppliers by account balance. On the structured version, NL2SQL methods (LLM-NL2SQL and XiYan) execute correctly by leveraging explicit fields such as “s_suppkey” and “s_acctbal”. On the semi-structured version, however, these attributes are embedded into free-text descriptions. The same models either output incorrect results or raise execution errors due to missing column references, illustrating the brittleness of NL2SQL pipelines under representational shifts. A similar failure is observed for the hybrid method Weaver, which also relies on SQL execution and cannot recover when key attributes are verbalized into text. This example highlights the challenges that semi-structured representations pose for models that depend on explicit schema linking and executable queries. We also provide additional case studies in Appendix F that highlight scenarios where different model families diverge in behavior.
@@ Appendix C Effect of Table Width
Appendix E Fine-grained Results
@@
Appendix F Additional Case Studies
@@
F.1 Model Behavior on Long Tables
@@
F.2 Model Behavior on Incomplete Schemas
@@
Appendix G Use of Large Language Models

### v3

@@ 4.3 Evaluation
Previous work (Deng et al., 2022; Wu et al., 2025a; Zhu et al., 2021) has largely relied on traditional metrics such as Exact Match (EM) and Partial Match (PM). However, these metrics suffer from well-known limitations: they depend on strict string-level comparison with the gold answer. Consequently, they penalize semantically correct predictions due to paraphrasing, formatting, or minor rounding differences. To address these limitations, we adopt LLMs as judges, a strategy shown to correlate strongly with human evaluation in a variety of reasoning tasks (Maekawa et al., 2025). Concretely, GPT-4o is given both the gold answer and the model’s prediction and prompted to determine semantic correctness. This evaluator tolerates surface variation while enforcing strict semantic equivalence. To assess the reliability of our LLM-based evaluator, we randomly sampled 100 evaluation cases across all datasets and model families. Each case was independently annotated by a human and compared to the LLM’s judgment. We observed a 96% agreement rate, demonstrating that the LLM-based evaluation is sufficiently stable for large-scale benchmarking. Unless stated otherwise, all results reported follow this evaluation protocol. We provide a Chi-Square significance analysis in Appendix E to rigorously evaluate performance differences across data splits.
@@ 5.6 Case Study
We now present illustrative examples that capture the main failure modes when shifting from structured to semi-structured representations. Figure 7 shows a representative case from the “supplier” table, where the query asks for the top ten suppliers by account balance. On the structured version, NL2SQL methods (LLM-NL2SQL and XiYan) execute correctly by leveraging explicit fields such as “s_suppkey” and “s_acctbal”. On the semi-structured version, however, these attributes are embedded into free-text descriptions. The same models either output incorrect results or raise execution errors due to missing column references, illustrating the brittleness of NL2SQL pipelines under representational shifts. A similar failure is observed for the hybrid method Weaver, which also relies on SQL execution and cannot recover when key attributes are verbalized into text. This example highlights the challenges that semi-structured representations pose for models that depend on explicit schema linking and executable queries. We also provide additional case studies in Appendix G that highlight scenarios where different model families diverge in behavior.
@@ Appendix C Effect of Table Width
Appendix E Statistical Significance Analysis
To rigorously evaluate the performance differences across various data splits, we conduct Chi-Square tests for each comparison. We report significance levels at $p<0.05$ (*), $p<0.01$ (**), and $p<0.001$ (***).
E.1 Table Length: Short vs. Long
Table 7 presents the performance comparison between short and long tables. All methods show statistically significant performance degradation when processing long tables ($p<0.001$), indicating that table length is a universal challenge. Gemini-2.5-flash exhibits the largest drop (52.17%), while LLM-NL2SQL and XiYan demonstrate relatively better robustness with drops of 11.87% and 11.06%, respectively.
Table 7: Performance comparison by table length (Short vs. Long).
Model Short (%) Long (%) Drop (%) p-value Sig.
GPT-4o 53.26 27.98 25.29 3.38e-21 ***
Gemini-2.5-flash 67.06 14.88 52.17 1.37e-83 ***
Qwen3-235B 46.72 24.24 22.48 2.32e-17 ***
LLM-NL2SQL 63.86 51.99 11.87 8.17e-06 ***
XiYan 54.06 43.00 11.06 4.75e-05 ***
H-STAR 53.59 12.19 41.40 3.70e-55 ***
Weaver 60.09 36.16 23.94 8.35e-19 ***
E.2 Number of Tables: Single vs. Multi
Table 8 shows the performance comparison between single-table and multi-table queries. Interestingly, not all methods are significantly affected by the number of tables. GPT-4o ($p=0.789$), Qwen3-235B ($p=0.097$), LLM-NL2SQL ($p=0.306$), and XiYan ($p=0.736$) show no statistically significant difference, suggesting robustness to table joins. In contrast, Gemini-2.5-flash and H-STAR perform significantly better on multi-table queries, while Weaver shows significant degradation.
Table 8: Performance comparison by number of tables (Single vs. Multi).
Model Single (%) Multi (%) $\Delta$ (%) p-value Sig.
GPT-4o 43.45 42.66 0.79 7.89e-01
Gemini-2.5-flash 46.81 59.03 -12.22 2.47e-06 ***
Qwen3-235B 38.00 33.81 4.19 9.71e-02
LLM-NL2SQL 59.25 56.57 2.68 3.06e-01
XiYan 49.77 50.76 -0.98 7.36e-01
H-STAR 37.75 57.34 -19.59 3.38e-14 ***
Weaver 52.89 33.39 19.50 4.97e-14 ***
E.3 Query Complexity: Lookup vs. Compositional
Table 9 presents the performance comparison between simple lookup queries and compositional reasoning queries. Most methods show significant performance degradation on compositional queries ($p<0.001$), with Qwen3-235B exhibiting the largest drop (37.79%). Notably, XiYan is the only method that shows no statistically significant difference ($p=0.060$), and even performs slightly better on compositional queries, suggesting its reasoning capability. LLM-NL2SQL also demonstrates relatively strong compositional reasoning with only a 9.75% drop.
Table 9: Performance comparison by query complexity (Lookup vs. Compositional).
Model Lookup (%) Comp. (%) Drop (%) p-value Sig.
GPT-4o 72.55 36.19 36.36 4.36e-33 ***
Gemini-2.5-flash 69.68 41.11 28.58 1.29e-20 ***
Qwen3-235B 68.24 30.45 37.79 9.77e-37 ***
LLM-NL2SQL 67.05 57.30 9.75 1.54e-03 **
XiYan 45.09 50.94 -5.85 6.01e-02
H-STAR 54.34 33.33 21.01 1.60e-12 ***
Weaver 73.98 45.02 28.96 3.10e-21 ***
E.4 Schema Quality: Clean vs. Noisy
Table 10 shows the performance comparison between clean and noisy (incomplete) schemas. All methods show statistically significant performance differences. LLM-NL2SQL shows the most severe degradation (54.62%), indicating high sensitivity to schema quality. Interestingly, H-STAR is the only method that performs better on noisy schemas ($p=0.035$), suggesting potential robustness to schema imperfections. Thanks to table preprocessing, Weaver maintains relatively stable performance with only an 8.64% drop.
Table 10: Performance comparison by schema quality (Clean vs. Noisy).
Model Clean (%) Noisy (%) Drop (%) p-value Sig.
GPT-4o 90.00 67.45 22.55 9.75e-17 ***
Gemini-2.5-flash 89.47 70.71 18.77 3.61e-12 ***
Qwen3-235B 87.37 62.13 25.24 1.37e-17 ***
LLM-NL2SQL 74.74 20.12 54.62 2.74e-47 ***
XiYan 52.64 24.56 28.08 1.40e-11 ***
H-STAR 77.90 84.91 -7.02 3.52e-02 *
Weaver 90.00 81.36 8.64 6.97e-04 ***
These results show that all the analysis supports our findings in the Section 5.
Appendix F Fine-grained Results
@@
Appendix G Additional Case Studies
@@
G.1 Model Behavior on Long Tables
@@
G.2 Model Behavior on Incomplete Schemas
@@
Appendix H Use of Large Language Models